Chemical Formula C6H12O 403 4,0i Opt 3.38 3.37 เก้ เท่ เก๋ เท่ เก้ เท่ เท่ เก่ เก่ ค 1.00 0.98 0.97 - --- -- 5.6 1 (ยาก) 6.0150 140 - 133.78 130 120 - 116.92 110 100 90 80 f1 (ppm) 70 - 73.55 - 70.02 60 50 40 30 - 22.93 20 10.62 10

The above is just one question. please answer the following according to the above.

1. Assign signals of 13C and 1H spectra as well as find the structural unit and functional group.

2. Give the structure

sample answer - as follows

1H NMR: p-substituted Ar at 7.27-7.97 ppm, benzylic CH3 group at 2.42 ppm, COOH proton at 12.9 ppm

13C NMR: carbonyl C at 167.8 ppm (ester or acid), benzylic CH3 at 21.40 ppm, 4 remaining signals are Ar carbons, two of which are quaternary, typical of a p-substituted ring

--> and finaly give a structure for the question.
Answer #1

Answer: Molecular formula = C6H120 Double bond equivalent = (C+1) - (H/2) = 7-(12/2) = (14-12)/2 = 1 The molecule contains 1From the HNMR data, the structure of the molecule is HNMR data(ppm) 0.98(3H, t) 1.60(2H, 9) 3.38(2H, 3) IH 4.02(2H,d) For-C13 C-NMR data 10.62(9) For-CH3 22.93(t) For-CH2 70.02(t) 73.55(t) For-CH2 bonded with oxygen atom For-CH2 bonded with oxygen